OpenStudy (anonymous):

(1 over 6)to the power of 4 times (1 over 6) to the power of negative 12... over (1 over 6) to the power of negative 3??????????????????????????????????/

jhonyy9 (jhonyy9):

\[(1/6)^{4} * (1/6)^{(-12)} / (1/6)^{(-3)}=(1/6)^{(4{+(-12)-(-3)}}\]

jhonyy9 (jhonyy9):

=(1/6) on the power of (-5)

jhonyy9 (jhonyy9):

this is understandably ?